OpenStudy (anonymous):

Please Help! Use the first three non-zero terms of the maclaurin series to approximate the value of the definite interval from 0 to 0.8 (1-cost)/t^2 d

OpenStudy (anonymous):

i could have sworn we did this. was it wrong?

OpenStudy (anonymous):

It was, I might of scrwed up the last part

OpenStudy (anonymous):

when we got to 1/2 - x^2/24 + x^4/720, I mustve screwed up after

OpenStudy (anonymous):

i recall the series was \[\frac{1}{2}-\frac{x^2}{24}+\frac{x^4}{720}\] plus stuff hold on let me check

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=mclauren+%281-cos%28x%29%29%2Fx^2 yup that part is right

OpenStudy (anonymous):

And then I probably screwed up with the integrating by term part

OpenStudy (anonymous):

technology gives me .39298 http://www.wolframalpha.com/input/?i=integral+0+to+.8+%281%2F2-x^2%2F24%2Bx^4%2F720%29

OpenStudy (anonymous):

might as well use it because you are not going to compute this with .8 without using a calculator, but the anti - derivative should be straightforward since you have a polynomial

OpenStudy (anonymous):

That's actually what ! got, when integrating by term I got sqrtx - x^3/450 + x^5/1440 .. is that good?

OpenStudy (anonymous):

However it also asks to do a test to see if it converges, and the ratio test doesn't work when I use that

OpenStudy (anonymous):

where id the root x come from?

OpenStudy (anonymous):

anti derivative of \(\frac{1}{2}\) is \(\frac{x}{2}\)

OpenStudy (anonymous):

Why do I have to take the anti-derivative if I can evaluate the integral before I take it?

OpenStudy (anonymous):

second question is also confusing. what does it mean "converge" when you only have three terms? unless you take the infinite series, you have a polynomial. the infinite series will of course converge but you will have write a formula for it

OpenStudy (anonymous):

you lost me on the last question your job is to find the definite integral right? a number

OpenStudy (anonymous):

I think I do have to write a formula for it, and I have to clue as to how to go from an integral to a series formula

OpenStudy (anonymous):

in other words \[\int_0^{.8}\frac{1}{2}-\frac{x^2}{24}+\frac{x^4}{720}dx\]

OpenStudy (anonymous):

to find this integral without using wolfram you need an anti derivative of \[\frac{1}{2}-\frac{x^2}{24}+\frac{x^4}{720}\]

OpenStudy (anonymous):

Yup that's right, kk I got that

OpenStudy (anonymous):

that is \[\frac{x}{2}-\frac{x^3}{72}+\frac{x^5}{3600}\] if my arithmetic is right

OpenStudy (anonymous):

the replace x by .8 and see what you get

OpenStudy (anonymous):

i gotta run, but post the question about convergence again, that is about convergence of the anti derivative. believe me it converges everywhere because you have a factorial in the denominator

OpenStudy (anonymous):

Thanks!!